# On the mixed UDA states and additivity

### v1

@@ III.5 On $n$-qubit mixed UDA states for $n\geq 5$
  • STEP 3 (Reduction of rank): $\rho_{1}$ is decomposed to the convex combination of some states $\alpha^{(i)}$, i.e. $\rho=\sum_{i}p_{i}\alpha^{(i)}$. The rank of $\alpha^{(i)}$ will be less than that of $\rho_{i}$ for most cases. The simplification of $\alpha^{(i)}$ is applied by LU equivalence. The set $B$ of states to be verified is constructed as follows. If $\alpha^{(i)}$ is a product state, i.e. $\alpha^{(i)}=\beta^{(i)}_{\cal K}\otimes\beta^{(i)}_{{\cal K}^{c}}$, then $B=\{\alpha^{(i)}\}\cup\{\beta^{(i)}_{\cal K},\beta^{(i)}_{{\cal K}^{c}}\}$; otherwise, $B=\{\alpha^{(i)}\}$. Using Lemma 4 and 5, if there is an element of $B$ that is not $k$-UDA by the necessary and sufficient conditions in Propositions 8 and 23, and Theorems 14, 20 and 22, we conclude that $\rho$ is not $k$-UDA; otherwise, we choose each element $B$ as $\rho$, and repeat STEP 1 to find possible states that is not $k$-UDA.
@@ IV Additivity of mixed UDA states
$\alpha_{A_{1}A_{2}}=|\psi\rangle\!\langle\psi|_{A_{j_{1}}}\otimes\alpha_{A_{j_{2}}},$ (16)
$\beta_{B_{1}B_{2}}=|\eta\rangle\!\langle\eta|_{B_{k_{1}}}\otimes\beta_{B_{k_{2}}},$ (17)
@@
$\alpha_{2}=|\psi_{2}\rangle\!\langle\psi_{2}|_{A_{2}}\otimes\alpha_{A_{1}}\otimes\alpha_{A_{3}}\;$ (20)
$(\text{or }\alpha_{2}=|\psi_{3}\rangle\!\langle\psi_{3}|_{A_{3}}\otimes\alpha_{A_{1}}\otimes\alpha_{A_{2}}),$ (21)
@@
$\beta_{2}=|\xi_{2}\rangle\!\langle\xi_{2}|_{B_{2}}\otimes\beta_{B_{1}}\otimes\alpha_{B_{3}}\;$ (25)
$(\text{or }\beta_{2}=|\xi_{3}\rangle\!\langle\xi_{3}|_{B_{3}}\otimes\beta_{B_{1}}\otimes\alpha_{B_{2}}),$ (26)
@@
with the positive diagonal elements of $D$ in (15). Suppose $\sigma$ is a state compatible with $\alpha_{j}\otimes\beta_{k}$ by all possible 2-marginals. It suffices to investigate the 2-UDA property of $\alpha_{j}\otimes\beta_{k}$. Without loss of generality, this derives ten cases technically, and the following analysis holds, up to swapping the $\alpha_{A_{j}}$’s with the $\beta_{B_{j}}$’s.
  1. 1.
    Suppose $\sigma$ is compatible with $\alpha_{1}\otimes\beta_{1}$. From Lemma 2, we have $\sigma=|\psi_{1}\rangle\!\langle\psi_{1}|_{A_{1}}\otimes|\xi_{1}\rangle\!\langle\xi_{1}|_{B_{1}}\otimes\sigma_{A_{2}A_{3}B_{2}B_{3}}$, where $\sigma_{A_{2}A_{3}}=\alpha_{A_{2}A_{3}}$, $\sigma_{B_{2}B_{3}}=\beta_{B_{2}B_{3}}$ and $\sigma_{A_{m}A_{n}}=\alpha_{A_{m}}\otimes\beta_{B_{n}}$. One can verify that $\alpha_{1}\otimes\beta_{1}$ is 2-UDA if and only if $\sigma_{A_{2}A_{3}B_{2}B_{3}}=\alpha_{A_{2}A_{3}}\otimes\beta_{B_{2}B_{3}}$. That is, $\alpha_{A_{2}A_{3}}\otimes\beta_{B_{2}B_{3}}$ is 2-UDA and thus satisfies the conditions in Theorem 20.
    Similarly, we consider the remaining nine cases and obtain the following facts.
  2. 2.
    $\alpha_{1}\otimes\beta_{2}$ is 2-UDA if and only if $\alpha_{A_{2}A_{3}}\otimes\beta_{B_{1}}\otimes\beta_{B_{3}}$ is 2-UDA. This is analyzed in Proposition 17.
  3. 3.
    $\alpha_{1}\otimes\beta_{3}$ is 2-UDA if and only if $\beta_{B_{1}}\otimes\alpha_{A_{2}A_{3}}$ is 2-UDA. That is, the resulting state $\beta_{B_{1}}\otimes\alpha_{A_{2}A_{3}}$ satisfies one of the conditions in Theorem 14.
  4. 4.
    $\alpha_{1}\otimes\beta_{4}$ is 2-UDA if and only if $\alpha_{A_{2}A_{3}}\otimes\beta_{B_{1}}\otimes(D\oplus 0)_{B_{2}B_{3}}$ is 2-UDA, which is analyzed in Theorem 22.
  5. 5.
    $\alpha_{2}\otimes\beta_{2}$ is 2-UDA if and only if $\alpha_{A_{1}}\otimes\alpha_{A_{3}}\otimes\beta_{B_{1}}\otimes\beta_{B_{3}}$ is 2-UDA. This is shown in Proposition 15.
  6. 6.
    $\alpha_{2}\otimes\beta_{3}$ is 2-UDA if and only if $\alpha_{A_{1}}\otimes\alpha_{A_{3}}\otimes\beta_{B_{1}}$ is 2-UDA. Using Proposition 9, $\alpha_{2}\otimes\beta_{3}$ is 2-UDA iff one of $\alpha_{A_{1}}$, $\alpha_{A_{3}}$ and $\beta_{B_{1}}$ is pure.
  7. 7.
    $\alpha_{2}\otimes\beta_{4}$ is 2-UDA iff $\alpha_{A_{1}}\otimes\alpha_{A_{3}}\otimes\beta_{B_{1}}\otimes(D\oplus 0)_{B_{2}B_{3}}$ is 2-UDA. This is considered in Proposition 23.
  8. 8.
    $\alpha_{3}\otimes\beta_{3}$ is 2-UDA without additional conditions.
  9. 9.
    $\alpha_{3}\otimes\beta_{4}$ is 2-UDA if and only if $\alpha_{A_{1}}\otimes\beta_{B_{1}}\otimes(D\oplus 0)_{B_{2}B_{3}}$ is 2-UDA by Theorem 20.
  10. 10.
    $\alpha_{4}\otimes\beta_{4}$ is 2-UDA iff $\alpha_{A_{1}}\otimes(H\oplus 0)_{A_{2}A_{3}}\otimes\beta_{B_{1}}\otimes(D\oplus 0)_{B_{2}B_{3}}$ is 2-UDA. From Theorem 22, this state is 2-UDA iff both $\alpha_{A_{1}}$ and $\beta_{B_{1}}$ are pure.
By the preceding analysis, we obtain the necessary and sufficient condition under which the additivity of two three-qubit UDA states holds.
@@
We recall that Theorem 1 in add2023shen shows that if two mixed states $\alpha$ and $\beta$ are both $k$-UDA and one of them is pure, then $\alpha\otimes\beta$ is $k$-UDA directly. Obviously, this fact is included in Theorem 27 (ii), where a pure $\alpha_{A_{j_{2}}A_{j_{3}}}\otimes\beta_{B_{1}}$ naturally satisfies the conditions in Theorem 14. It is also shown that the purity of $\alpha$ or $\beta$ is a rather strict condition for mixed states since the additivity also holds for mixed UDA $\alpha$ and $\beta$. Theorem 27 implies that the rank of $\alpha$ and $\beta$ is less than four and the rank of $\alpha\otimes\beta$ is less than 4 if their UDA additivity holds. Note that $\alpha\otimes\beta$ is a six-qubit state whose rank is at most 64. Hence, UDA additivity holds only for states with relatively low rank.
@@ V Applications
The reconstruction of $\rho$ usually requires the expectation values $\pi_{i_{1}\cdots i_{n}}$ of $D_{U}=4^{n}-1$ Pauli observables and many samples of $\rho$. The complexity is later lowered in fully quantum state tomography (FQST), where $D_{F}=3^{n}$ global Pauli observables are required by ignoring $\sigma_{0}$ in each qubit. However, FQST is unfeasible for large systems. By using the complete characterization of UDA states, global states can be reconstructed from measurements on $k$-marginals, a technique known as local measurements on reduced density matrices (LQST). Generally, LQST requires measuring $D_{L}=3^{k}\binom{n}{k}$ local observables. The Parallel quantum state tomography (PQST) protocol for $n$-qubit states experiment2024hu exploits parallel Pauli measurements to efficiently obtain all $k$-marginals, and reconstructs the global state via tensor-network learning. By choosing $k=2$, PQST requires $D_{P}^{(2)}=3+6\lceil\log_{2}n\rceil$ observables.
@@
Three-qubit state $\rho_{A_{1}A_{2}A_{3}}$ in Theorem 20 63 27 27 1515
Four-qubit state $\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$ in Theorem 20 255 81 54 1515
@@
  • Entanglement detection and distribution. When an entangled state is UDA, entanglement can be certified indirectly through local measurements yu2020multi; shi2025entangle. Otherwise, a non-$k$-UDA state implies the irreducible $(k+1)$-partite correlations irr2008zhou.
  • Security in quantum communication and cryptography. An adversary may attempt to distribute a counterfeit global state that reproduces the same local statistics as a desired resource state. If the target state is UDA with respect to the monitored data, such forgery is impossible huang2025certify; swingle2014reconstruct.
  • Quantum error correction and recoverability. A UDA state reflects a form of information redundancy: the global quantum state is uniquely encoded in partial subsystems. This is related to quantum error correction, where recovering a state from partial information is a key requirement gupta2024encode.
  • Characterization of many-body quantum phases. States exhibiting topological order or long-range entanglement are generally not UDA with respect to local reduced states lee2025topo, whereas short-range entangled phases often admit UDA properties. This connects UDA states to the study of quantum phases and phase classification will2025probing.
@@ VI Conclusion
Many problems arising from this work can be further explored. First, we know that a full-rank state is not UDA, while a low-rank state may be UDA. It is interesting to analyze the threshold of the rank of a mixed state, where states whose rank exceeds this threshold are not UDA. Second, the necessary and sufficient condition by which arbitrary multipartite states can be further considered. The techniques of purification of mixed states and range inclusion may be utilized. Then the UDA additivity of arbitrary multipartite mixed states can be further investigated.
@@ Appendix A Proof of preliminary lemmas
We denote by $A_{[m]}$ and $B_{[n]}$ the $m$-partite system $(A_{1},...,A_{m})$ and $n$-partite system $(B_{1},...,B_{n})$, respectively. Let $\rho_{A[m]B[n]}=\alpha_{A_{[m]}}\otimes\beta_{B_{[n]}}$. If $\rho$ is $k$-UDA, then for any state $\sigma$ compatible with $\rho$, i.e. $\rho_{{\cal S}}=\sigma_{{\cal S}}$ for ${\cal S}\subset A_{[m]}B_{[n]}$ and $|{\cal S}|=k$, it holds that $\rho=\sigma$. Since $\alpha$ and $\beta$ are both not $k$-UDA, there are states $\eta_{A_{[m]}}$ and $\xi_{B_{[n]}}$ such that $\eta_{{\cal S}_{a}}=\alpha_{{\cal S}_{a}}$ for ${\cal S}_{a}\subset A_{[m]}$ and $\xi_{{\cal S}_{b}}=\beta_{{\cal S}_{b}}$ for ${\cal S}_{b}\subset B_{[m]}$, and $\alpha\neq\eta$, $\beta\neq\xi$. We claim that $\eta\otimes\xi$ has the same $k$-marginals as that of $\alpha\otimes\beta$. If ${\cal S}\subset A_{[m]}$ or ${\cal S}\subset B_{[n]}$, the claim holds by choosing ${\cal S}={\cal S}_{a}$ or ${\cal S}={\cal S}_{b}$. Otherwise, suppose ${\cal S}=(S_{1},S_{2})$ and $S_{1}\in A_{[m]}$, $S_{2}\in B_{[n]}$. Let ${\cal G}_{1}=\{S_{1},S_{l}\}\subset A_{[m]}$ and ${\cal G}_{2}=\{S_{2},S_{j}\}\subset B_{[n]}$. Then
$\hskip-8.5359pt(\eta\otimes\xi)_{{\cal S}}=$ $\mathop{\rm Tr}_{\{A_{[m]}\backslash S_{1}\}\cup\{B_{[n]}\backslash S_{2}\}}(\eta\otimes\xi)$
$=$ $\mathop{\rm Tr}_{S_{k}}\big[\mathop{\rm Tr}_{\{A_{[m]}\backslash{\cal G}_{1}\}}(\eta)\big]\otimes\mathop{\rm Tr}_{S_{j}}\big[\mathop{\rm Tr}_{\{B_{[n]}\backslash{\cal G}_{2}\}}(\xi)\big]$
$=$ $\mathop{\rm Tr}_{S_{k}}(\alpha_{{\cal G}_{1}})\otimes\mathop{\rm Tr}_{S_{j}}(\beta_{{\cal G}_{2}})=\alpha_{S_{1}}\otimes\beta_{S_{2}}$
$=$ $(\alpha\otimes\beta)_{{\cal S}}=\rho_{\cal S}.$
We choose $\sigma=\eta\otimes\xi$. Then $\sigma_{{\cal S}}=\rho_{{\cal S}}$ for any ${\cal S}\subset A_{[m]}B_{[n]}$. Since $\rho$ is $k$-UDA, we have $\rho=\sigma$. However, $\alpha\neq\eta$ and $\beta\neq\xi$ implies that $\alpha\otimes\beta\neq\eta\otimes\xi$. In fact, we assume that $\alpha=(a_{ij})$ and $\beta=(b_{ij})$ be two normalized quantum states. From $\alpha\neq\beta$, there is $a_{ij}\neq b_{ij}$ and without loss of generality, $a_{ij}\neq 0$. Suppose $b_{ij}=ha_{ij}$ for $h\neq 1$. If $\alpha\otimes\beta=\eta\otimes\xi$, then $a_{ij}\beta=b_{ij}\xi$. So $\beta=h\xi$ and $\beta$ is not normalized. Hence we have $\rho\neq\sigma$. It is a contradiction.   $\sqcap\sqcup$
@@
Let $\alpha$ and $\beta$ be two states of systems $(A_{1}A_{2}...A_{m})$ and $(B_{1}B_{2}...B_{n})$, respectively. Suppose $\alpha\otimes\beta$ is not 2-UDA. There is a different $(m+n)$-partite state $\sigma_{A_{[m]}B_{[n]}}$ that is $k$-compatible with $\alpha\otimes\beta$, i.e. $\sigma_{A_{\cal K}B_{\cal P}}=\alpha_{A_{\cal K}}\otimes\beta_{B_{\cal P}}$ with $|A_{\cal K}|+|B_{\cal P}|=k$. Obviously, $\sigma_{B_{\cal P}A_{\cal K}}=\beta_{B_{\cal P}}\otimes\alpha_{A_{\cal K}}$. That is, the $k$-partite marginals of the state $\sigma_{B_{[n]}A_{[m]}}$ is the same as that of $\beta\otimes\alpha$. Since $\sigma_{A_{[m]}B_{[n]}}\neq\alpha\otimes\beta$, we have $\sigma_{B_{[n]}A_{[m]}}\neq\beta\otimes\alpha$. This completes the proof.   $\sqcap\sqcup$

### v2

@@ III.5 On $n$-qubit mixed UDA states for $n\geq 5$
  • STEP 3 (Reduction of rank): $\rho_{1}$ is decomposed to the convex combination of some states $\alpha^{(i)}$, i.e. $\rho=\sum_{i}p_{i}\alpha^{(i)}$. The rank of $\alpha^{(i)}$ will be less than that of $\rho_{i}$ for most cases. The simplification of $\alpha^{(i)}$ is applied by LU equivalence. The set $B$ of states to be verified is constructed as follows. If $\alpha^{(i)}$ is a product state, i.e. $\alpha^{(i)}=\beta^{(i)}_{\cal K}\otimes\beta^{(i)}_{{\cal K}^{c}}$, then $B=\{\alpha^{(i)}\}\cup\{\beta^{(i)}_{\cal K},\beta^{(i)}_{{\cal K}^{c}}\}$; otherwise, $B=\{\alpha^{(i)}\}$. Using Lemma 4 and 5, if there is an element of $B$ that is not $k$-UDA by the necessary and sufficient conditions in Propositions 9 and 25, and Theorems 15, 22 and 24, we conclude that $\rho$ is not $k$-UDA; otherwise, we choose each element $B$ as $\rho$, and repeat STEP 1 to find possible states that is not $k$-UDA.
@@ IV Additivity of mixed UDA states
$\alpha_{A_{1}A_{2}}=|\psi\rangle\!\langle\psi|_{A_{j_{1}}}\otimes\alpha_{A_{j_{2}}},\quad\beta_{B_{1}B_{2}}=|\eta\rangle\!\langle\eta|_{B_{k_{1}}}\otimes\beta_{B_{k_{2}}},$ (16)
@@
$\alpha_{2}=|\psi_{2}\rangle\!\langle\psi_{2}|_{A_{2}}\otimes\alpha_{A_{1}}\otimes\alpha_{A_{3}}\;(\text{or }\alpha_{2}=|\psi_{3}\rangle\!\langle\psi_{3}|_{A_{3}}\otimes\alpha_{A_{1}}\otimes\alpha_{A_{2}}),$ (19)
@@
$\beta_{2}=|\xi_{2}\rangle\!\langle\xi_{2}|_{B_{2}}\otimes\beta_{B_{1}}\otimes\alpha_{B_{3}}\;(\text{or }\beta_{2}=|\xi_{3}\rangle\!\langle\xi_{3}|_{B_{3}}\otimes\beta_{B_{1}}\otimes\alpha_{B_{2}}),$ (23)
@@
with the positive diagonal elements of $D$ in (15). Suppose $\sigma$ is a state compatible with $\alpha_{j}\otimes\beta_{k}$ by all possible 2-marginals. It suffices to investigate the 2-UDA property of $\alpha_{j}\otimes\beta_{k}$. Without loss of generality, this derives ten cases technically, and the following analysis holds, up to swapping the $\alpha_{A_{j}}$’s with the $\beta_{B_{j}}$’s. Suppose $\sigma$ is compatible with $\alpha_{1}\otimes\beta_{1}$. From Lemma 2, we have $\sigma=|\psi_{1}\rangle\!\langle\psi_{1}|_{A_{1}}\otimes|\xi_{1}\rangle\!\langle\xi_{1}|_{B_{1}}\otimes\sigma_{A_{2}A_{3}B_{2}B_{3}}$, where $\sigma_{A_{2}A_{3}}=\alpha_{A_{2}A_{3}}$, $\sigma_{B_{2}B_{3}}=\beta_{B_{2}B_{3}}$ and $\sigma_{A_{m}A_{n}}=\alpha_{A_{m}}\otimes\beta_{B_{n}}$. One can verify that $\alpha_{1}\otimes\beta_{1}$ is 2-UDA if and only if $\sigma_{A_{2}A_{3}B_{2}B_{3}}=\alpha_{A_{2}A_{3}}\otimes\beta_{B_{2}B_{3}}$. That is, $\alpha_{A_{2}A_{3}}\otimes\beta_{B_{2}B_{3}}$ is 2-UDA and thus satisfies the conditions in Theorem 22. Similarly, we consider the remaining nine cases and obtain the necessary and sufficient condition under which the additivity of two three-qubit UDA states holds.
@@
We recall that Theorem 1 in Shen and Chen (2023) shows that if two mixed states $\alpha$ and $\beta$ are both $k$-UDA and one of them is pure, then $\alpha\otimes\beta$ is $k$-UDA directly. Obviously, this fact is included in Theorem 29 (ii), where a pure $\alpha_{A_{j_{2}}A_{j_{3}}}\otimes\beta_{B_{1}}$ naturally satisfies the conditions in Theorem 15. It is also shown that the purity of $\alpha$ or $\beta$ is a rather strict condition for mixed states since the additivity also holds for mixed UDA $\alpha$ and $\beta$. Theorem 29 implies that the rank of $\alpha$ and $\beta$ is less than four and the rank of $\alpha\otimes\beta$ is less than 4 if their UDA additivity holds. Note that $\alpha\otimes\beta$ is a six-qubit state whose rank is at most 64. Hence, UDA additivity holds only for states with relatively low rank.
@@ V Applications
The reconstruction of $\rho$ usually requires the expectation values $\pi_{i_{1}\cdots i_{n}}$ of $D_{U}=4^{n}-1$ Pauli observables and many samples of $\rho$. The complexity is later lowered in fully quantum state tomography (FQST), where $D_{F}=3^{n}$ global Pauli observables are required by ignoring $\sigma_{0}$ in each qubit. However, FQST is unfeasible for large systems. By using the complete characterization of UDA states, global states can be reconstructed from measurements on $k$-marginals, a technique known as local measurements on reduced density matrices (LQST). Generally, LQST requires measuring $D_{L}=3^{k}\binom{n}{k}$ local observables. The Parallel quantum state tomography (PQST) protocol for $n$-qubit states Hu et al. (2024) exploits parallel Pauli measurements to efficiently obtain all $k$-marginals, and reconstructs the global state via tensor-network learning. By choosing $k=2$, PQST requires $D_{P}^{(2)}=3+6\lceil\log_{2}n\rceil$ observables.
@@
Three-qubit state $\rho_{A_{1}A_{2}A_{3}}$ in Theorem 22 63 27 27 $15$
Four-qubit state $\rho_{A_{1}A_{2}}\otimes\rho_{A_{3}A_{4}}$ in Theorem 22 255 81 54 $15$
@@
  • Entanglement detection and distribution. When an entangled state is UDA, entanglement can be certified indirectly through local measurements Yu (2020); Shi et al. (2025). Otherwise, a non-$k$-UDA state implies the irreducible $(k+1)$-partite correlations Zhou (2008).
  • Security in quantum communication and cryptography. An adversary may attempt to distribute a counterfeit global state that reproduces the same local statistics as a desired resource state. If the target state is UDA with respect to the monitored data, such forgery is impossible Huang et al. (2025a); Swingle and Kim (2014).
  • Quantum error correction and recoverability. A UDA state reflects a form of information redundancy: the global quantum state is uniquely encoded in partial subsystems. This is related to quantum error correction, where recovering a state from partial information is a key requirement Gupta et al. (2024).
  • Characterization of many-body quantum phases. States exhibiting topological order or long-range entanglement are generally not UDA with respect to local reduced states Lee and Moon (2025), whereas short-range entangled phases often admit UDA properties. This connects UDA states to the study of quantum phases and phase classification Will et al. (2025).
@@ VI Conclusion
Many problems arising from this work can be further explored. First, we know that a full-rank state is not UDA, while a low-rank state may be UDA. It is interesting to analyze the threshold of the rank of a mixed state, where states whose rank exceeds this threshold are not UDA. Second, the necessary and sufficient condition by which arbitrary multipartite states can be further considered. The techniques of purification of mixed states and range inclusion may be utilized. Besides, merging two systems as a local one may help to the determination of UDA states. For instance, the four-partite state $\alpha:=\alpha_{A}\otimes\alpha_{BCD}$ is 2-UDA may imply that a tripartite state $\beta$ is 2-UDA, where $\beta$ is the same as $\alpha$, but regarded as a tripartite state by merging systems $A$ and $B$. The UDA additivity of arbitrary multipartite mixed states can be further investigated.
@@ Appendix A Proof of preliminary lemmas
We denote by $A_{[m]}$ and $B_{[n]}$ the $m$-partite system $(A_{1},...,A_{m})$ and $n$-partite system $(B_{1},...,B_{n})$, respectively. Suppose $\alpha_{A_{[m]}}$ is not $k$-UDA and $\beta_{B_{[n]}}$ is arbitrary. Then there is another state $\sigma_{A_{[m]}}$ that is $k$-compatible with $\alpha$, i.e. $\alpha_{\cal S}=\sigma_{\cal S}$ for $|{\cal S}|=k$. We show that $\sigma\otimes\beta$ is $k$-compatible with $\alpha\otimes\beta$. Let ${\cal S}=(S_{1},S_{2})$ and $S_{1}\in A_{[m]}$, $S_{2}\in B_{[n]}$.
$(\sigma\otimes\beta)_{\cal S}=\mathop{\rm Tr}_{\{A_{[m]}\backslash S_{1}\}\cup\{B_{[n]}\backslash S_{2}\}}(\sigma\otimes\beta)=\sigma_{S_{1}}\otimes\beta_{S_{2}}=\alpha_{S_{1}}\otimes\beta_{S_{2}}=(\alpha\otimes\beta)_{\cal S},$
where $\alpha_{S_{1}}=\sigma_{S_{1}}$ by $S_{1}\subseteq{\cal S}$. From $\alpha\neq\sigma$, we have $\alpha\otimes\beta\neq\sigma\otimes\beta$. Hence, $\alpha\otimes\beta$ is not $k$-UDA.   $\sqcap\sqcup$
@@
(i) Let $\alpha$ and $\beta$ be two states of systems $(A_{1}A_{2}...A_{m})$ and $(B_{1}B_{2}...B_{n})$, respectively. Suppose $\alpha\otimes\beta$ is not 2-UDA. There is a different $(m+n)$-partite state $\sigma_{A_{[m]}B_{[n]}}$ that is $k$-compatible with $\alpha\otimes\beta$, i.e. $\sigma_{A_{\cal K}B_{\cal P}}=\alpha_{A_{\cal K}}\otimes\beta_{B_{\cal P}}$ with $|A_{\cal K}|+|B_{\cal P}|=k$. Obviously, $\sigma_{B_{\cal P}A_{\cal K}}=\beta_{B_{\cal P}}\otimes\alpha_{A_{\cal K}}$. That is, the $k$-partite marginals of the state $\sigma_{B_{[n]}A_{[m]}}$ is the same as that of $\beta\otimes\alpha$. Since $\sigma_{A_{[m]}B_{[n]}}\neq\alpha\otimes\beta$, we have $\sigma_{B_{[n]}A_{[m]}}\neq\beta\otimes\alpha$.   $\sqcap\sqcup$
Lemma 8 shows the general properties of UDA states. The proof is shown here.
Proof.
(i) Let ${\cal S}\subset(A_{1}A_{2}...A_{n})$ and $|{\cal S}|=k$. For $\rho=\text{Re}(\rho)+i\text{Im}(\rho)$, we have $\text{Re}(\rho)_{\cal S}=\frac{1}{2}(\rho_{\cal S}+\rho^{*}_{\cal S})$. Since $\rho$ is $k$-compatible with $\sigma$, we have $\rho_{\cal S}=\sigma_{\cal S}$ and $\rho^{*}_{\cal S}=\sigma^{*}_{\cal S}$. Then $\text{Re}(\rho)_{\cal S}=\frac{1}{2}(\sigma_{\cal S}+\sigma^{*}_{\cal S})=\text{Re}(\sigma)_{\cal S}$. Hence $\rho$ has the same real parts as $\sigma$. The equivalence of imaginary parts $\text{Im}(\rho)$ and $\text{Im}(\sigma)$ can be shown similarly.
(ii) Suppose $A:=\text{Re}(\rho)=\frac{\rho+\rho^{*}}{2}$ and $B:=\text{Im}(\rho)=\frac{\rho-\rho^{*}}{2i}$, which are symmetric and antisymmetric, respectively. Let $z=x+iy\in\mathbb{C}^{n}$, for $x,y\in\mathbb{R}^{n}$. We have $z^{\dagger}\rho z=\begin{bmatrix}x^{T}&y^{T}\end{bmatrix}M\begin{bmatrix}x\\ y\end{bmatrix}\geq 0$, for $M=\begin{bmatrix}A&B\\ -B&A\end{bmatrix}$. That is, $M\geq 0$. For any $u=x+iy\in\ker(A)$, we have $Ax=Ay=0$. By choosing $v_{1}=(x,0)^{T}$ and $v_{2}=(0,y)^{T}$, we have $Bx=By=0$, i.e. $u\in\ker(B)$. Then ${\cal R}(B^{\dagger})\subseteq{\cal R}(A^{\dagger})$, i.e. ${\cal R}(B)\subseteq{\cal R}(A)$.
(iii) Let $\rho=\text{Re}(\rho)+i\text{Im}(\rho)$, for $\text{Im}(\rho)\neq 0$. We have $\rho^{*}\neq\rho$. Since all $k$-marginals of $\rho$ are real, we have $\rho_{\cal S}=\rho_{{\cal S}}^{*}$ with $|{\cal S}|=k$. That is, $\rho^{*}$ is another state $k$-compatible with $\rho$.
(iv) It suffices to show that $\ker(\rho_{\cal S})\subseteq\ker(H_{\cal S})$. For any $|x\rangle\in\ker(\rho_{\cal S})$, we have $\rho_{\cal S}|x\rangle=0$, which is equivalent to $\langle x|\rho_{\cal S}|x\rangle=0$. Since $\langle x|\rho_{\cal S}|x\rangle=\sum_{j}\langle x,j|\rho|x,j\rangle=0$, we have $\rho|x,j\rangle=0$, i.e. $|x,j\rangle\in\ker(\rho)$. By $\ker(\rho)\subseteq\ker(H)$, we have $|x,j\rangle\in\ker(H)$. We show that $|x\rangle\in\ker H_{\cal S}$. For any $|u\rangle\in{\cal H}_{\cal S}$, we have $\langle x|H_{S}|u\rangle=\sum_{j}\langle x,j|H|u,j\rangle=0$. Hence $H_{\cal S}|x\rangle=0$, i.e. $|x\rangle\in\ker(H_{\cal S})$. The converse does not generally hold, for example, $H=|01\rangle\!\langle 01|$ and $\rho=|\psi\rangle\!\langle\psi|$ with $|\psi\rangle=|00\rangle+|11\rangle$.
(v) Let ${\cal S}\subset[A_{1}A_{2}...A_{n}]$ and $|{\cal S}|=k$. If $A_{1}\notin{\cal S}$, we have $(\Lambda_{A_{1}}\otimes I)\eta]_{\cal S}=\eta_{\cal S}$ for $\eta=\alpha,\;\beta$. By $\alpha_{\cal S}=\beta_{\cal S}$, we prove the claim. Otherwise, we have $A_{1}\in{\cal S}$ and $(\Lambda_{A_{1}}\otimes I)\eta]_{\cal S}=(\Lambda_{A_{1}}\otimes I_{{\cal S}\backslash A_{1}})\eta_{\cal S}$. Hence $(\Lambda_{A_{1}}\otimes I)\alpha]_{\cal S}=[(\Lambda_{A_{1}}\otimes I)\beta]_{\cal S}$, and $(\Lambda_{A_{1}}\otimes I_{B})\alpha$ and $(\Lambda_{A_{1}}\otimes I_{B})\beta$ are $k$-compatible. Further, if $\Lambda_{A_{1}}$ is reversible, there is another channel $R_{A}$, such that $(R_{A}\circ\Lambda_{A})\otimes I]\alpha=\alpha$. If $(\Lambda_{A}\otimes I)\alpha$ is not $k$-UDA, then there is another state $\beta$ $k$-compatible with it. By the preceding analysis, $(R_{A}\otimes I)\beta$ is $k$-compatible with $\beta$, and thus compatible with $\alpha$. Here $(R_{A}\otimes I)\beta\neq\alpha$ by $(\Lambda_{A}\otimes I)\alpha\neq\beta$, which derives that $\alpha$ is not $k$-UDA. Similarly, $\alpha$ is not $k$-UDA derives that $(\Lambda_{A}\otimes I)\alpha$ is not $k$-UDA.   $\sqcap\sqcup$